# Proper scoring rules for estimation and forecast evaluation

### v1

@@ 1 Introduction
A probabilistic forecast is a probability distribution over the space of the possible outcomes of the future event that is stated by the forecaster. The simplest and most popular case of probabilistic forecasts arises when the outcome is binary, so the probabilistic forecast reduces to issuing a predictive probability of success. Brier (1950) was the first to consider the problem of devising a scoring rule which could not be “played” by a dishonest forecasting agent. He introduced the quadratic scoring rule and showed that it incentivizes a forecasting agent to state his most accurate probability estimate when faced with uncertainty. McCarthy (1956)111better known as the designer of the Lisp programming language. was the first to clearly state the concept of propriety as it is understood today and stated at (1).
@@
The literature knows both orientations in that a proper scoring rule $S(\mathbb{P},\mathbb{Q})$ should be minimized at $\mathbb{P}=\mathbb{Q}$ or that it should be maximized. The influential paper by Gneiting and Raftery (2007) assumes that better forecasts correspond to higher scoring rule values, but most of the more recent literature has adopted the convention (1). With a negative orientation, scoring rules should maybe be called loss functions, but it is not in our power to change the naming convention. “Scoring rule” is sometimes abbreviated to “score”, which may cause confusion, since “score” typically refers to the gradient of the logarithmic density in statistics, machine learning and econometrics. However, for some classes of proper scoring rules such as kernel scores (Section 3.1) the term is established by now.
@@ 1.1 Proper scoring rules in estimation
If we think of $\mathbb{Q}$ as the data distribution and $\mathbb{P}$ as a model distribution, then statement (1) can be interpreted as saying that the expectation $S(\cdot,\mathbb{Q})=\mathbb{E}_{Y\sim\mathbb{Q}}[S(\cdot,Y)]$ is minimized precisely when the model and data distributions are equal. It is reasonable to expect that given an iid sample $\smash{\{Y_{j}\}_{j=1}^{n}}$ from $\mathbb{Q}$ with empirical distribution ^$\hat{\mathbb{Q}}_{n}$, the minimizer of the empirical mean $S(\cdot,\hat{\mathbb{Q}}_{n})=(1/n)\sum_{j=1}^{n}S(\cdot,Y_{j})$ would give reasonably good estimates of $\mathbb{Q}$ under appropriate regularity conditions. Scoring rule minimization can thus be employed for estimating probability distributions.
@@
where $S$ is a strictly proper scoring rule. This is known as minimum scoring rule inference, of which the classical maximum likelihood inference is a special case where $S=S_{\mathrm{log}}$ is the logarithmic score of Example 2 (Dawid et al., 2016). Strict propreity ensures that if $\hat{\theta}$ was minimizing the limit of $S(\mathbb{P}_{\theta},\hat{\mathbb{Q}}_{n})$ instead, which is the expectation $S(\mathbb{P}_{\theta},\mathbb{P}_{\theta^{\ast}})=S(\mathbb{P}_{\theta},\mathbb{Q})$, then $\hat{\theta}=\theta_{\ast}$. Propriety can thus be thought of as the population counterpart of consistency for M-estimators of probability distributions (Huber and Ronchetti, 2009).
@@ 1.2 Proper scoring rules in forecast evaluation
Clearly, the empirical scores at (3) can be also be used as an evaluation metric (validation or testing loss) for comparing the performance of different estimates.
While (1) ensures that the optimal forecaster (or true model) will dominate all other forecasters (models or parameters) with regards to the comparison at (3) in the long term, different proper scoring rules will rank competing imperfect forecasters differently even for infinite amounts of data, see for example Patton (2020). While this may seem as a disadvantage of ranking forecasts with respect to proper scoring rules, it is neither surprising nor is there any alternative. Therefore, the scoring rule must be chosen carefully for the comparison of empirical scores to be relevant to the application at hand, since it determines when one imperfect probabilistic forecast should be preferred over another. As a result, there are numerous proposals to tailor proper scoring rules such that they are discriminative with respect to certain regions of interest of the predictive distributions and outcomes, see Section 5.
@@ 2.1 Definition
for every $\mathbb{P},\mathbb{Q}\in\mathcal{P}$ and strictly proper if (4) holds with equality if and only if $\mathbb{P}=\mathbb{Q}$.
@@
Let $\mu$ be a $\sigma$-finite measure on $\mathcal{Y}$ and $\mathcal{P}$ the set of measures which are absolutely continuous with respect to $\mu$. For example, if $\mathcal{Y}=\mathbb{R}^{d}$ and $\mu$ is the Lebesgue measure, we consider absolutely continuous distributions on $\mathbb{R}^{d}$; if $\mathcal{Y}$ is finite and $\mu$ is the counting measure, we have categorical random variables. The logarithmic score $S_{\mathrm{log}}:\mathcal{P}\times\mathcal{Y}\to\overline{\mathbb{R}}$ with $S_{\mathrm{log}}(\mathbb{P},y)=-\log p(y)$ is a strictly proper scoring rule (Good, 1952). The logarithmic score is an example of a local scoring rule, see Section 3.2 for further details. Minimizing the logarithmic score is equivalent to the well-known maximum likelihood principle, which is the most popular method for estimating probability distributions and arguably, the foundation of classical statistical inference.
@@
Let $\mathcal{Y}=\mathbb{R}^{d}$ and $\mathcal{P}$ be the set of absolutely continuous distributions with twice continuously differentiable densities $p$ such that $\|\nabla_{\mathbf{x}}\log p(\mathbf{x})\|\to 0$ as $\|\mathbf{x}\|\to\infty$. The Hyvärinen score $S_{\mathrm{Hyv\ddot{a}rinen}}:\mathcal{P}\times\mathcal{Y}\to\overline{\mathbb{R}}$ given by
@@
where $p$ is the density of $\mathbb{P}$, is a strictly proper scoring rule (Parry et al., 2012). Minimizing the Hyvärinen score is equivalent to score matching, which is the basis of the score-based modeling techniques behind many of the recent advances in machine learning.
@@
Let $\mathcal{Y}=\mathbb{R}$ and $\mathcal{P}$ be the set of distributions with finite mean. The continuous ranked probability score $\operatorname{CRPS}:\mathcal{P}\to\overline{\mathbb{R}}$ given by $\operatorname{CRPS}(\mathbb{P},y)=\int\left(F_{\mathbb{P}}(x)-\mathbf{1}\{y\leq x\}\right)^{2}\,\mathrm{d}x$, where $F_{\mathbb{P}}$ is the CDF of $\mathbb{P}$, is a strictly proper scoring rule (Matheson and Winkler, 1976). It has compelling interpretation as the $L^{2}$ distance between the CDFs of $\mathbb{P}$ and the forecast $y$. Alternatively, it can also be seen as the Brier score for the binary event $\mathbf{1}{\{Y\leq x\}}$ integrated over $\mathbb{R}$. Furthermore, has the equivalent representation
@@
see for example Baringhaus and Franz (2004, Lemma 2.1). The representation at (5) identifies the CRPS as a kernel score, see Section 3.1. It is of great significance in meteorology, where it is the most popular scoring rule for continuous univariate quantities.
@@
With the goal of forecast evaluation in mind, the notion of propriety at (4) is easily motivated, compare (3). However, the notion of propriety also appears naturally from the perspective of (distributional) regression, compare (2). The following result has previously appeared in Tsyplakov (2011); Holzmann and Eulert (2014). It is analogous to a result of Banerjee et al. (2005) on conditional expectation and Bregman loss functions.
@@
Mathematically, the concept of proper scoring rules is an extension to probability distributions of the concept of Bregman loss functions (Ovcharov, 2018). These are distance-like functions on Euclidean spaces used in the optimization literature to adapt general methods such as proximal gradient descent to the geometry of specific problems (Teboulle, 2018). Importantly, many boosting algorithms such as AdaBoost are known to be instances of one such method known as mirror descent (Collins et al., 2002; Beja-Battais, 2023).
@@
$H(\mathbb{P})=\int S(\mathbb{P},y)\,\mathrm{d}\mathbb{P}(y)\qquad\mbox{ and }\qquad d(\mathbb{P},\mathbb{Q})=S(\mathbb{P},\mathbb{Q})-H(\mathbb{Q}).$ (6)
@@
$H(\mathbb{P})=\tfrac{1}{2}\int\|\nabla\log p(\mathbf{x})\|^{2}\,p(\mathbf{x})\,\mathrm{d}\mathbf{x}.$
The associated divergence $d(\mathbb{P},\mathbb{Q})=\tfrac{1}{2}\int\|\nabla\log p(\mathbf{y})-\nabla\log q(\mathbf{y})\|^{2}\,p(\mathbf{y})\,\mathrm{d}\mathbf{y}$ is sometimes called score matching distance.
@@ 2.2 Characterization
At (7), we use the shorthand $\langle f,\mathbb{P}\rangle$ for the integral $\int f\,\mathrm{d}\mathbb{P}$ of a $\mathbb{P}$-integrable function $f$ on $\mathcal{Y}$, and $\delta_{y}$ denotes the Dirac measure at the point $y$. Furthermore, a supergradient of a concave function $H\colon\mathcal{P}\to\overline{\mathbb{R}}$ at $\mathbb{P}$ is a function $h_{\mathbb{P}}\colon\mathcal{Y}\to\overline{\mathbb{R}}$ that is $\mathbb{Q}$-integrable and
@@
for all $\mathbb{Q}\in\mathcal{P}$. The function $H$ at (7) indeed coincides with the entropy of $S$ defined at (6). A regular proper scoring rule is thus determined by its entropy $H$ and the family $\{h_{\mathbb{P}}:\mathbb{P}\in\mathcal{P}\}$ of chosen supergradients. Theorem 12 is general but also somewhat abstract since supergradients of concave functions on convex sets of probability measures may not be directly intuitive. However, we will showcase the strength and direct applicability of this result in several instances.
@@
Typically, the entropy $H$ is differentiable the sense that for every $\mathbb{P}\in\mathcal{P}$, there exists a $\mathcal{P}$-integrable function $\nabla_{\mathbb{P}}H:\mathcal{Y}\to\overline{\mathbb{R}}$ such that
@@ 2.3 Geometric Properties
Let $H:\mathcal{P}\to\overline{\mathbb{R}}$ be a twice-differentiable function and let $d:\mathcal{P}\times\mathcal{P}\to\overline{\mathbb{R}}$ denote the divergence of the corresponding scoring rule. Then
@@
Let $\mathbb{P},\mathbb{Q}$ be absolutely continuous distributions on $\mathbb{R}^{d}$ with Lebesgue densities $p$ and $q$, respectively. If $\mathbb{P}$ and $\mathbb{Q}$ are “close”, the divergence $d(\mathbb{P},\mathbb{Q})$ with respect to the logarithmic score, quadratic score (see (16) in Section 3.3), and CRPS is approximately given by the right-hand side of (10), which is, respectively,
@@ 3.1.1 Construction and computation
for some conditionally negative definite kernel $h:\mathcal{Y}\times\mathcal{Y}\to[0,\infty)$, and have first been discussed by Dawid (2007); Gneiting and Raftery (2007). A function $h:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$ is a conditionally negative definite kernel, if it is symmetric, that is, $h(x,y)=h(y,x)$ for $x,y\in\mathcal{Y}$, and for every $n\geq 1$, $x_{1},\dots,x_{n}\in\mathcal{Y}$ and $\alpha_{1},\dots,\alpha_{n}\in\mathbb{R}$ with $\sum_{j=1}^{n}\alpha_{j}=0$, it satisfies $\sum_{i,j=1}^{n}\alpha_{i}\alpha_{j}h(x_{i},x_{j})\leq 0$. Moreover, we assume that $h(x,y)\geq 0$ for $x,y\in\mathcal{Y}$. This does not lead to any loss of generality, see Remark 25 in Appendix A.
@@
For any conditionally negative definite kernel $h$, the entropy at (11) is concave, and hence the induced scoring rule is proper. Strict propriety is a more delicate issue. Assume $h\geq 0$ and let $\mathcal{P}_{h}$ be the set of all probability measures such that $H(\mathbb{P})<\infty$. The kernel $h$ is strongly conditionally negative definite if $\iint h(x,y)\,\mathrm{d}(\mathbb{P}-\mathbb{Q})(x)\,\mathrm{d}(\mathbb{P}-\mathbb{Q})(y)=0$ implies $\mathbb{P}=\mathbb{Q}$ for any $\mathbb{P},\mathbb{Q}\in\mathcal{P}_{h}$. Strongly conditionally negative definite kernels are closely related to positive definite characteristic and universal kernels, see Gretton et al. (2006); Sriperumbudur et al. (2011).
@@
A kernel score is a (strictly) proper scoring rule $S:\mathcal{P}_{h}\times\mathcal{Y}\to\overline{\mathbb{R}}$ of the form (12) for some (strongly) conditionally negative definite kernel $h:\mathcal{Y}\times\mathcal{Y}\to[0,\infty)$. The divergence $d$ of a kernel score is $d(\mathbb{P},\mathbb{Q})=-\frac{1}{2}\iint h(y,y^{\prime})\,\mathrm{d}(\mathbb{P}-\mathbb{Q})(y)\,\mathrm{d}(\mathbb{P}-\mathbb{Q})(y^{\prime}).$ We can write $S(\mathbb{P},y)=d(\mathbb{P},\delta_{y})+\tfrac{1}{2}h(y,y)$ since $\delta_{y}\in\mathcal{P}_{h}$ for $y\in\mathcal{Y}$. In fact, $(\mathbb{P},y)\mapsto d(\mathbb{P},y)$ is itself a proper scoring rule and is strongly equivalent to $S$. The intimate connection between conditionally negative definite kernels and the geometry of Hilbert spaces results in many interesting interpretations of the kernel score that are best stated in terms of its divergence.
@@
The kernel score $S:\mathcal{P}\times\mathcal{Y}\to\overline{\mathbb{R}}$ as in (12) is a (strictly) proper scoring rule if and only if there exists a Hilbert space $\mathcal{H}$ and a subset $\{\psi_{x}\}_{x\in\mathcal{Y}}\subseteq\mathcal{H}$ such that the divergence $d$ of $S$ satisfies
@@
We can thus construct a kernel score for a given outcome space $\mathcal{Y}$ as soon as we can construct a conditionally negative definite kernel on $\mathcal{Y}$ or a Hilbertian embedding of $\mathcal{Y}$. A simple approach is to choose $h=-k$ for some strictly positive definite kernel $k:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$ which can be constructed in many ways (Williams and Rasmussen, 2006). For example, $k(\mathbf{x},\mathbf{y})=-\exp(-\|\mathbf{x}-\mathbf{y}\|^{p}/\lambda)$ corresponds to the Laplacian ($p=1$) and Gaussian ($p=2$) kernels on $\mathbb{R}^{d}$ leading to bounded kernel scores. Kernel scores implicitly correspond to different embeddings of the outcome space $\mathcal{Y}$ into a Hilbert space via (13). For example, $(\mathbf{x},\mathbf{y})\mapsto\|\mathbf{x}-\mathbf{y}\|^{\beta/2}$ for $\beta\in(0,2)$ is actually a metric on $\mathcal{Y}=\mathbb{R}^{d}$. According to a well-known result of Schoenberg (1937), the corresponding metric space can be embedded into a Hilbert space, which implies the existence of an embedding $\{\psi_{\mathbf{x}}\}_{\mathbf{x}\in\mathcal{Y}}\subseteq\mathcal{H}$ such that $\|\psi_{\mathbf{x}}-\psi_{\mathbf{y}}\|_{\mathcal{H}}=\|\mathbf{x}-\mathbf{y}\|^{\beta/2}$ for $\mathbf{x},\mathbf{y}\in\mathcal{Y}$.
@@
$\hat{S}(\mathbb{P},y)=\frac{1}{2m(m-1)}\sum_{i=1}^{m}\sum_{i^{\prime}:i^{\prime}\neq i}h(X_{i},X_{i^{\prime}})-\frac{1}{m}\sum_{i=1}^{m}h(X_{i},y)$ (14)

### v2

@@ 1 Introduction
A probabilistic forecast is a probability distribution over the space of the possible outcomes of the future event that is stated by the forecaster. The simplest and most popular case of probabilistic forecasts arises when the outcome is binary, so the probabilistic forecast reduces to issuing a predictive probability of success. Brier (1950) was the first to consider the problem of devising a scoring rule which could not be “played” by a dishonest forecasting agent. He introduced the quadratic scoring rule and showed that it incentivizes a forecasting agent to state his most accurate probability estimate when faced with uncertainty. McCarthy (1956)111better known as the designer of the Lisp programming language. was the first to clearly state the concept of propriety as it is understood today and stated as Equation 1.
@@
The literature knows both orientations in that a proper scoring rule $S(\mathbb{P},\mathbb{Q})$ should be minimized at $\mathbb{P}=\mathbb{Q}$ or that it should be maximized. The influential paper by Gneiting and Raftery (2007) assumes that better forecasts correspond to higher scoring rule values, but most of the more recent literature has adopted the convention in Equation 1. With a negative orientation, scoring rules should maybe be called loss functions, but it is not in our power to change the naming convention.
Some parts of the literature order the arguments of proper scoring rules differently, that is, it would be $S(y,\mathbb{P})$ instead of $S(\mathbb{P},y)$. We mention this since it may cause confusion when dealing with expected scores $S(\mathbb{P},\mathbb{Q})$. “Scoring rule” is sometimes abbreviated to “score”, which may cause confusion, since “score” typically refers to the gradient of the logarithmic density in statistics, machine learning and econometrics. However, for some classes of proper scoring rules such as kernel scores (Section 3.1) the term is established by now.
@@ 1.1 Proper scoring rules in estimation
If we think of $\mathbb{Q}$ as the data distribution and $\mathbb{P}$ as a model distribution, then Equation 1 can be interpreted as saying that the expectation $S(\cdot,\mathbb{Q})=\mathbb{E}_{Y\sim\mathbb{Q}}[S(\cdot,Y)]$ is minimized precisely when the model and data distributions are equal. It is reasonable to expect that given an iid sample $\smash{\{Y_{j}\}_{j=1}^{n}}$ from $\mathbb{Q}$ with empirical distribution ^$\hat{\mathbb{Q}}_{n}$, the minimizer of the empirical mean $S(\cdot,\hat{\mathbb{Q}}_{n})=(1/n)\sum_{j=1}^{n}S(\cdot,Y_{j})$ would give reasonably good estimates of $\mathbb{Q}$ under appropriate regularity conditions. Scoring rule minimization can thus be employed for estimating probability distributions.
@@
where $S$ is a strictly proper scoring rule. This is known as minimum scoring rule inference, of which the classical maximum likelihood inference is a special case where $S=S_{\mathrm{log}}$ is the logarithmic score of Example 2 (Dawid et al., 2016). Strict propriety ensures that if $\hat{\theta}$ was minimizing the limit of $S(\mathbb{P}_{\theta},\hat{\mathbb{Q}}_{n})$ instead, which is the expectation $S(\mathbb{P}_{\theta},\mathbb{P}_{\theta^{\ast}})=S(\mathbb{P}_{\theta},\mathbb{Q})$, then $\hat{\theta}=\theta_{\ast}$. Propriety can thus be thought of as the population counterpart of consistency for M-estimators of probability distributions (Huber and Ronchetti, 2009).
@@ 1.2 Proper scoring rules in forecast evaluation
Clearly, the empirical scores in Equation 3 can be also be used as an evaluation metric (validation or testing loss) for comparing the performance of different estimates.
While Equation 1 ensures that the optimal forecaster (or true model) will dominate all other forecasters (models or parameters) with regards to the comparison in Equation 3 in the long term, different proper scoring rules will rank competing imperfect forecasters differently even for infinite amounts of data, see for example Patton (2020). While this may seem as a disadvantage of ranking forecasts with respect to proper scoring rules, it is neither surprising nor is there any alternative. Therefore, the scoring rule must be chosen carefully for the comparison of empirical scores to be relevant to the application at hand, since it determines when one imperfect probabilistic forecast should be preferred over another. As a result, there are numerous proposals to tailor proper scoring rules such that they are discriminative with respect to certain regions of interest of the predictive distributions and outcomes, see Section 5.
@@ 2.1 Definition
for every $\mathbb{P},\mathbb{Q}\in\mathcal{P}$ and strictly proper if Equation 4 holds with equality if and only if $\mathbb{P}=\mathbb{Q}$.
@@
Let $\mu$ be a $\sigma$-finite measure on $\mathcal{Y}$ and $\mathcal{P}$ the set of measures which are absolutely continuous with respect to $\mu$. For example, if $\mathcal{Y}=\mathbb{R}^{d}$ and $\mu$ is the Lebesgue measure, we consider absolutely continuous distributions on $\mathbb{R}^{d}$; if $\mathcal{Y}$ is finite and $\mu$ is the counting measure, we have categorical random variables. The logarithmic score $S_{\mathrm{log}}:\mathcal{P}\times\mathcal{Y}\to\overline{\mathbb{R}}$ with $S_{\mathrm{log}}(\mathbb{P},y)=-\log p(y)$, where $p$ is the density of $\mathbb{P}$ with respect to $\mu$, is a strictly proper scoring rule (Good, 1952). The logarithmic score is an example of a local scoring rule, see Section 3.2 for further details. Minimizing the logarithmic score is equivalent to the well-known maximum likelihood principle, which is the most popular method for estimating probability distributions and arguably, the foundation of classical statistical inference.
@@
[…]
@@
where $\nabla_{\mathbf{y}}f$ and $\Delta_{\mathbf{y}}f$ denote the gradient $(\partial f/\partial y_{1},\dots,\partial f/\partial y_{d})$ and Laplacian $\sum_{j=1}^{d}\partial^{2}f/\partial y_{j}^{2}$ of $f:\mathbb{R}^{d}\mapsto\mathbb{R}$ with respect to $\mathbf{y}=(y_{1},\dots,y_{d})$, and $p$ is the density of $\mathbb{P}$, is a strictly proper scoring rule (Parry et al., 2012). Minimizing the Hyvärinen score is equivalent to score matching, which is the basis of the score-based modeling techniques behind many of the recent advances in machine learning.
@@
Let $\mathcal{Y}=\mathbb{R}$ and $\mathcal{P}$ be the set of distributions with finite mean. The continuous ranked probability score $\operatorname{CRPS}:\mathcal{P}\to\overline{\mathbb{R}}$ given by $\operatorname{CRPS}(\mathbb{P},y)=\int\left(F_{\mathbb{P}}(x)-\mathbf{1}\{y\leq x\}\right)^{2}\,\mathrm{d}x$, where $F_{\mathbb{P}}$ is the CDF of $\mathbb{P}$, is a strictly proper scoring rule (Matheson and Winkler, 1976). It has compelling interpretation as the $L^{2}$ distance between the CDFs of $\mathbb{P}$ and the forecast $y$. Alternatively, it can also be seen as the Brier score for the binary event $\mathbf{1}{\{Y\leq x\}}$ integrated over $\mathbb{R}$. Furthermore, it has the equivalent representation
@@
see for example Baringhaus and Franz (2004, Lemma 2.1). The representation in Equation 5 identifies the CRPS as a kernel score, see Section 3.1. It is of great significance in meteorology, where it is the most popular scoring rule for real-valued variables.
@@
With the goal of forecast evaluation in mind, the notion of propriety at Equation 4 is easily motivated, compare Equation 3. However, the notion of propriety also appears naturally from the perspective of (distributional) regression, compare Equation 2. The following result has previously appeared in Tsyplakov (2011); Holzmann and Eulert (2014). It is analogous to a result of Banerjee et al. (2005) on conditional expectation and Bregman loss functions.
@@
Mathematically, the concept of proper scoring rules is an extension of the concept of Bregman loss functions to probability distributions (Ovcharov, 2018). These are distance-like functions on Euclidean spaces used in the optimization literature to adapt general methods such as proximal gradient descent to the geometry of specific problems (Teboulle, 2018). Importantly, many boosting algorithms such as AdaBoost are known to be instances of one such method known as mirror descent (Collins et al., 2002; Beja-Battais, 2023).
@@
$H(\mathbb{P})=S(\mathbb{P},\mathbb{P})=\int S(\mathbb{P},y)\,\mathrm{d}\mathbb{P}(y)\qquad\mbox{ and }\qquad d(\mathbb{P},\mathbb{Q})=S(\mathbb{P},\mathbb{Q})-H(\mathbb{Q}).$ (6)
@@
$H(\mathbb{P})=-\tfrac{1}{2}\int\|\nabla\log p(\mathbf{x})\|^{2}\,p(\mathbf{x})\,\mathrm{d}\mathbf{x},$
see Remark 1 in the appendix for derivation. The associated divergence $d(\mathbb{P},\mathbb{Q})=\tfrac{1}{2}\int\|\nabla\log p(\mathbf{y})-\nabla\log q(\mathbf{y})\|^{2}\,q(\mathbf{y})\,\mathrm{d}\mathbf{y}$ is sometimes called score matching distance.
@@ 2.2 Characterization
In Equation 7, we use the shorthand $\langle f,\mathbb{P}\rangle$ for the integral $\int f\,\mathrm{d}\mathbb{P}$ of a $\mathbb{P}$-integrable function $f$ on $\mathcal{Y}$, and $\delta_{y}$ denotes the Dirac measure at the point $y$. Furthermore, a supergradient of a concave function $H\colon\mathcal{P}\to\overline{\mathbb{R}}$ at $\mathbb{P}$ is a function $h_{\mathbb{P}}\colon\mathcal{Y}\to\overline{\mathbb{R}}$ that is $\mathbb{Q}$-integrable and
@@
for all $\mathbb{Q}\in\mathcal{P}$. The function $H$ in Equation 7 indeed coincides with the entropy of $S$ defined in Equation 6. A regular proper scoring rule is thus determined by its entropy $H$ and the family $\{h_{\mathbb{P}}:\mathbb{P}\in\mathcal{P}\}$ of chosen supergradients. Theorem 12 is general but also somewhat abstract since supergradients of concave functions on convex sets of probability measures may not be directly intuitive. However, we will showcase the strength and direct applicability of this result in several instances.
@@
Typically, the entropy $H$ is differentiable in the sense that for every $\mathbb{P}\in\mathcal{P}$, there exists a $\mathcal{P}$-integrable function $\nabla_{\mathbb{P}}H:\mathcal{Y}\to\overline{\mathbb{R}}$ such that
@@ 2.3 Geometric Properties
Let $H:\mathcal{P}\to\overline{\mathbb{R}}$ be a twice-differentiable concave function and let $d:\mathcal{P}\times\mathcal{P}\to\overline{\mathbb{R}}$ denote the divergence of the corresponding scoring rule. Then
@@
Let $\mathbb{P},\mathbb{Q}$ be absolutely continuous distributions on $\mathbb{R}^{d}$ with Lebesgue densities $p$ and $q$, respectively. If $\mathbb{P}$ and $\mathbb{Q}$ are “close”, the divergence $d(\mathbb{P},\mathbb{Q})$ with respect to the logarithmic score, quadratic score (see Equation 16 in Section 3.3), and CRPS is approximately given by the right-hand side of Equation 10, which is, respectively,
@@ 3.1.1 Construction and computation
for some conditionally negative definite kernel $h:\mathcal{Y}\times\mathcal{Y}\to[0,\infty)$, and have been popularized by Dawid (2007); Gneiting and Raftery (2007). A function $h:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$ is a conditionally negative definite kernel, if it is symmetric, that is, $h(x,y)=h(y,x)$ for $x,y\in\mathcal{Y}$, and for every $n\geq 1$, $x_{1},\dots,x_{n}\in\mathcal{Y}$ and $\alpha_{1},\dots,\alpha_{n}\in\mathbb{R}$ with $\sum_{j=1}^{n}\alpha_{j}=0$, it satisfies $\sum_{i,j=1}^{n}\alpha_{i}\alpha_{j}h(x_{i},x_{j})\leq 0$. Moreover, we assume that $h(x,y)\geq 0$ for $x,y\in\mathcal{Y}$. This does not lead to any loss of generality, see Remark 2 in the appendix.
@@
For any conditionally negative definite kernel $h$, the entropy at Equation 11 is concave, and hence the induced scoring rule is proper. Strict propriety is a more delicate issue. Assume $h\geq 0$ and let $\mathcal{P}_{h}$ be the set of all probability measures such that $H(\mathbb{P})<\infty$. The kernel $h$ is strongly conditionally negative definite if $\iint h(x,y)\,\mathrm{d}(\mathbb{P}-\mathbb{Q})(x)\,\mathrm{d}(\mathbb{P}-\mathbb{Q})(y)=0$ implies $\mathbb{P}=\mathbb{Q}$ for any $\mathbb{P},\mathbb{Q}\in\mathcal{P}_{h}$. Strongly conditionally negative definite kernels are closely related to positive definite characteristic and universal kernels, see Gretton et al. (2006); Sriperumbudur et al. (2011).
@@
A kernel score is a (strictly) proper scoring rule $S:\mathcal{P}_{h}\times\mathcal{Y}\to\overline{\mathbb{R}}$ of the form in Equation 12 for some (strongly) conditionally negative definite kernel $h:\mathcal{Y}\times\mathcal{Y}\to[0,\infty)$. The divergence $d$ of a kernel score is $d(\mathbb{P},\mathbb{Q})=-\frac{1}{2}\iint h(y,y^{\prime})\,\mathrm{d}(\mathbb{P}-\mathbb{Q})(y)\,\mathrm{d}(\mathbb{P}-\mathbb{Q})(y^{\prime}).$ We can write $S(\mathbb{P},y)=d(\mathbb{P},\delta_{y})+\tfrac{1}{2}h(y,y)$ since $\delta_{y}\in\mathcal{P}_{h}$ for $y\in\mathcal{Y}$. In fact, $(\mathbb{P},y)\mapsto d(\mathbb{P},y)$ is itself a proper scoring rule and is strongly equivalent to $S$. The intimate connection between conditionally negative definite kernels and the geometry of Hilbert spaces results in many interesting interpretations of the kernel score that are best stated in terms of its divergence.
@@
The function $S:\mathcal{P}\times\mathcal{Y}\to\overline{\mathbb{R}}$ as in Equation 12 is a (strictly) proper scoring rule if and only if there exists a Hilbert space $\mathcal{H}$ and a subset $\{\psi_{x}\}_{x\in\mathcal{Y}}\subseteq\mathcal{H}$ such that the divergence $d$ of $S$ satisfies
@@
We can thus construct a kernel score for a given outcome space $\mathcal{Y}$ as soon as we can construct a conditionally negative definite kernel on $\mathcal{Y}$ or a Hilbertian embedding of $\mathcal{Y}$. A simple approach is to choose $h=-k$ for some strictly positive definite kernel $k:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$ which can be constructed in many ways (Williams and Rasmussen, 2006). For example, $k(\mathbf{x},\mathbf{y})=-\exp(-\|\mathbf{x}-\mathbf{y}\|^{p}/\lambda)$ corresponds to the Laplacian ($p=1$) and Gaussian ($p=2$) kernels on $\mathbb{R}^{d}$ leading to bounded kernel scores. Kernel scores implicitly correspond to different embeddings of the outcome space $\mathcal{Y}$ into a Hilbert space via Equation 13. For example, $(\mathbf{x},\mathbf{y})\mapsto\|\mathbf{x}-\mathbf{y}\|^{\beta/2}$ for $\beta\in(0,2)$ is actually a metric on $\mathcal{Y}=\mathbb{R}^{d}$. According to a well-known result of Schoenberg (1937), the corresponding metric space can be embedded into a Hilbert space, which implies the existence of an embedding $\{\psi_{\mathbf{x}}\}_{\mathbf{x}\in\mathcal{Y}}\subseteq\mathcal{H}$ such that $\|\psi_{\mathbf{x}}-\psi_{\mathbf{y}}\|_{\mathcal{H}}=\|\mathbf{x}-\mathbf{y}\|^{\beta/2}$ for $\mathbf{x},\mathbf{y}\in\mathcal{Y}$.
@@
$\hat{S}(\mathbb{P},y)=\frac{1}{m}\sum_{i=1}^{m}h(X_{i},y)-\frac{1}{2m(m-1)}\sum_{i=1}^{m}\sum_{i^{\prime}:i^{\prime}\neq i}h(X_{i},X_{i^{\prime}})$ (14)